# Ubiquity of the spin-orbit induced magnon nonreciprocity in ultrathin ferromagnets

### v1

@@ Abstract
The propagation of magnons along a symmetry path may depend on the direction of propagation, similar to many other quasiparticles in nature. This phenomenon is commonly referred to as nonreciprocity. In addition to the fact that it is of great interest to understand the fundamental physical mechanism leading to this nonreciprocal propagation, the phenomenon of magnon nonreciprocity may be used to design magnon-based logic devices. Recently, it has been demonstrated that a significantly large spin-orbit coupling can lead to giant nonreciprocity of exchange-dominated terahertz magnons, when they are excited by means of spin-polarized electrons [Phys. Rev. Lett. 132, 126702 (2024)]. Here, by providing experimental results on two additional systems we demonstrate the generality of the observed phenomenon. Comparing the results of a Co/Ni bilayer on Ir(001) to those of a Co double layer on Ir(001) and W(110) we unravel the impact of the interfacial electronic hybridization on the observed phenomenon and provide further insights into the microscopic mechanism leading to this nonreciprocal magnon excitation. It was observed that the interfacial electronic hybridization, which is governing the magnetic properties of the interface atomic layer, is somewhat important but is not crucial for the magnon nonreciprocity. On the other hand the choice of the incident energy of the incoming electron beam is decisive for the observation of the effect. Our results indicate that depending on the energy of the incident electron beam and the scattering geometry the magnon nonreciprocity can be tuned and even be inverted for some ranges of the magnon momentum. The results may inspire new ideas for the generation of nonreciprocal ultrafast terahertz magnons in ultrathin films of $3d$ magnetic metals as well as in two-dimensional van der Waals magnets, interfaced with solids possessing a large spin-orbit coupling, for applications in the field of magnonics.
@@ I Introduction
Usually when a quasiparticle is propagating along two opposite but equivalent directions it is expected to exhibit identical properties, e.g., amplitude, lifetime, propagation speed, etc. [1, 2]. However, under some circumstances these properties may depend on the direction of propagation. This phenomenon is usually referred to as nonreciprocity and has been expected to be observed for magnons –the quanta of spin waves in magnetic materials– when they are propagating along two opposite directions of a symmetry path. The effect is of great importance from both fundamental as well as application points of view. From a fundamental point of view, a nonreciprocal propagation of a quasiparticle is associated with the presence of an interaction, which possibly breaks the time-reversal symmetry. It is, therefore, interesting to unravel the nature of the nonreciprocity, in order to address the presence of such a time-reversal symmetry breaking interaction. From an application point of view, nonreciprocal propagation of magnons is of great importance to realize novel magnon-based logic devices useful for applications in the field of magnonics and spintronics, as has already been demonstrated in the case of magnons with gigahertz frequencies [3, 4, 5, 6, 7, 8, 9].
@@
Recently, it has been demonstrated that, indeed, such a competition between the spin-orbit and exchange scattering mechanisms during the scattering event of spin-polarized electrons from ferromagnetic surfaces possessing a large spin-orbit coupling (SOC) leads to a nonreciprocal excitation of high-energy terahertz (THz) magnons [15].
It is worth emphasizing that the magnon nonreciprocity is expected to be observed for the exchange-dominated THz magnons. Owing to the ultrafast propagation of these type of magnons, they would allow for information processing on ultrashort time scales [16]. Hence, such magnons are of great interest to the field of ultrafast magnonics and spintronics [17].
@@
In order to unravel the microscopic origin of the spin-orbit induced magnon nonreciprocity we investigated the impact of two important factors on the magnon nonreciprocity: (ii) the interfacial electronic hybridization, and (ii) the atomic structure and the symmetry of the of the magnetic overlayer as well as the choice of the substrate.
@@
Although the effect is expected to be observed by other excitation schemes, it should be best investigated by means of spin-polarized electron scattering experiments, e.g., spin-polarized high-resolution electron energy-loss spectroscopy (SPHREELS). While the original experiments were performed on a Co double layer on Ir(001), in order to investigate the impact of the interfacial electronic hybridization on the magnon nonreciprocity, we replaced the interface Co atomic layer with a Ni layer. We note that both Co double layer and Co/Ni bilayer possess nearly the same geometrical structure (crystallographic structure and the interatomic distances) [18, 19, 20]. Since most of the $3d$ states of Ni are occupied and are located below the Fermi level, it is expected that these states hybridize rather weakly with that of the $5d$ states of the Ir substrate [21, 22]. This would provide a unique possibility to examine the influence of the interfacial electronic hybridization on the observed magnon nonreciprocity. On the other hand considering the available $5d$ states of the substrate, the tungsten exhibits a lower number of occupied $5d$ states compared to iridium. Hence, it would be interesting to replace the Ir(001) substrate with W(110) and see how it changes the magnon nonreciprocity. Unlike Ir, which possesses a face-centered cubic structure, the crystal structure of W is body-centered cubic. Since in the case of Co double layer on W(110) the atomic arrangement is also different, one would, therefore, expect to observe effects associated with the change in the atomic arrangement as well. In order to address the above mentioned points we investigated the details of the spin-dependent electron scattering from two systems, i.e., Co/Ni/Ir(001) and 2Co/W(110) and compare the results to those of 2Co/Ir(001), investigated previously [15]. We will show that, although the interfacial electronic hybridization is crucial for the determination the magnetic properties of the magnetic layers grown on a nonmagnetic substrate, they are not decisive for the observed magnon nonreciprocity. In contrary, the magnon nonreciprocity depends strongly on the incident electron energy and the scattering geometry. Moreover, for some ranges of magnon wavevectors the magnon nonreciprocity can even be inverted.
@@
The paper is organized as following. In Sec. II we provide the experimental details. The substrate preparation, film growth and the SPHREELS experiments are explained. The results of the two systems, namely Co/Ni/Ir(001) and 2Co/W(110) are presented in Secs. III.1 and III.2, respectively, and are discussed in detail. A concluding remark with the perspective of the observation of the effect in other layered magnetic structures is provides in Sec. IV.
@@ II.1 Substrate preparation
In this study we used Ir(001) and W(110) as substrates. The choice of Ir(001) is based on the fact that one can grow well-ordered ultrathin films of $3d$ ferromagnetic metals on this substrate. In particular considering Co and Ni as overlayers the growth mode, lattice structure and the interatomic distances are very similar [18, 23, 24]. They both grow in a face-centered tetragonal structure [18, 23, 24]. This is particularly important to address the effects associated with the changes in the chemical properties of the interface layer without changing its geometrical structures. The choice of W(110) was based on the fact that it also exhibits a rather large SOC. Moreover, one can grow high-quality ultrathin films of $3d$ ferromagnets on this surface [25, 26, 27, 28, 29].
@@
Figure 1: The magneto-optical Kerr effect (MOKE) hysteresis loops recorded on (a) Co/Ni/Ir(001) and (b) 2Co/W(110) samples. The loops were recorded in longitudinal geometry at room temperature. In both cases the external magnetic field was applied along the [$\bar{1}$0]-direction of the substrates. The red and blue colors indicate the upward and downward magnetic field sweep, respectively.
@@ II.3 Spin-polarized high-resolution electron energy-loss spectroscopy
It is well-known that the magnon excitation process in SPHREELS is based on the exchange process [38, 39, 40, 33]. Since a magnon carries a total angular momentum of $1\hbar$ it can only be excited by incidence of a minority electron (when $\bm{P}$ is parallel to $\bm{M}$). This means that in fact the difference spectra $I_{\mu\downarrow}-I_{\mu\uparrow}$, which describe a spin-flip process, shall contain all the information regarding the magnons as spin-flip excitations. Since the magnon excitation is mediated by the exchange process, in the absence of SOC the excitation amplitude should be independent from the direction of $\bm{M}$ with respect to the crystallographic directions. It should only depend on the relative orientation of $\bm{P}$ and $\bm{M}$. Hence, the intensity of the difference spectra is expected to be equal for the two magnetization directions, despite a negative sign. The presence of SOC during the scattering event leads to an inequality of the intensity of the difference spectra, since it couples the spin to the crystallographic directions. On the other hand reversing the direction of $\bm{M}$ is equivalent to reversing the direction of the magnon wavevector [41, 42, 29]. This is because in a ferromagnet the time-reversal symmetry $\mathcal{T}$ is naturally broken. Hence, reversing the direction of $\bm{M}$ is equivalent to a time-reversal experiment. Therefore, the magnons for two opposite magnetization directions shall represent cases in which the magnons propagate along opposite directions. It has been shown that the presence of SOC leads to the fact that the magnon excitation amplitude, as seen by SPHREELS, is no longer equal for the two opposite directions of $\bm{M}$. The phenomenon is understood in terms of a strong competition between spin-orbit and exchange scattering mechanisms [15]. While in most of ferromagnetic elements such as Fe, Co and Ni SOC is not significantly large, the spin-orbit scattering is expected to be small. However, in ultrathin films of ferromagnetic metals grown on substrates with a large SOC spin-orbit scattering can be significant. The presence of a large SOC can lead to an inequality of magnon amplitudes, a phenomenon known as magnon nonreciprocity. In order to understand the effect in detail we prepared and investigated two additional systems as will be discussed in the following section.
@@ III.1 The case of Co/Ni/Ir(001)
The observed magnon nonreciprocity is also present for magnons having other wavevectors (see the discussion below). The difference spectra presented in Fig. 3 contain all the information regarding the excited magnon wavepackets, e.g., amplitude, lifetime, frequency, group and phase velocity, etc. One may, therefore, construct the magnon wavepackets propagating along two opposite directions based on these information [43, 44]. Since the excitation energy (the magnon eigenfrequency), the lifetime, group and phase velocities are nearly the same for the two wavepackets, we realized that the main factor, which leads to the nonreciprocal propagation of the magnon wavepackets is due the inequality of the amplitude of the wavepackets in the initial stage of propagation (just at the time when they are excited). The amplitude of the magnon wavepackets is directly given by the height of the difference spectra at the magnon excitation energy. Hence, it would be very useful to analyze, in detail, the dependence of the ratio of the peak heights of magnon excitation spectra for $\bm{M}\parallel[1\bar{1}0]$ and $\bm{M}\parallel[\bar{1}10]$ [see Fig. 3(c)]. This quantity may also be called as the magnon nonreciprocity. For a reciprocal propagation one expects a value equal to unity and for a unidirectional propagation this value should be zero. Since the absolute value of the intensity of the excitation spectra depends on the scattering geometry one should use the normalized difference spectra [similar to those shown in Fig. 3(c)]. In Fig. 4 the excitation peak intensities for the two magnetization directions as well as their ratio are presented for different values of $Q$. One immediate conclusion from the data presented in Fig. 4 is that the magnon nonreciprocity depends strongly on $Q$ and can even be inverted for some ranges of $Q$.
@@
Figure 5: Spin asymmetries of elastically scattered electrons against the incident electron energy $E_{i}$. $A_{\bm{M}\parallel[1\bar{1}0]}$ and $A_{\bm{M}\parallel[\bar{1}10]}$ were recorded for two opposite directions of magnetization, $[1\bar{1}0]$ and $[\bar{1}10]$, respectively, as described in the text. Different contributions to the spin asymmetries are $A_{SO}$ and $A_{Ex}$ and are given by Eqs. (1) and (2), respectively.
@@
Data shown in Fig. 5 indicate that both SOA and ExA exhibit a strong dependence on $E_{i}$. The behavior is very similar to that of the Co double layer on the same substrate [15]. In particular, the pattern of SOA is very similar in both systems, indicating that the SOC is mainly supported by the Ir(001) substrate. Since the interfacial electronic hybridization is expected to be marginal in the case of Ni/Ir interface, another important conclusion from this observation is that such a hybridization does not influence the behavior of SOA, drastically. In order to carefully analyze the spin asymmetries one may consider several different regions of $E_{i}$. For instance, below 7.5 eV, $A_{SO}$ and $A_{Ex}$ possess a similar magnitude but an opposite sign (shown by shaded yellow color in Fig. 5). In the range of 7.5 eV$<E_{i}<11.5$ eV $A_{SO}$ and $A_{Ex}$ have the same sign but a different magnitude (light-blue shaded area in Fig. 5). For the incident energies above 11.5 eV $A_{SO}$ and $A_{Ex}$ exhibit again an opposite sign, similar to the first region (yellow shaded area in Fig. 5). In order to assess the consequences of the incident energy on the magnon nonreciprocity we analyzed the behavior of these asymmetries as a function of $Q$ for two areas marked by A and B in Fig. 5. Note that one could also analyze the data for $E_{i}<7.5$ eV. The results are expected to be similar to those of area B (see the data provided for the 2Co/W(110) system in Sec. III.2). Both SOA and ExA were measured as a function of $Q$ for two different values of the incident energy (i) $E_{i}=10$ eV, located in area A, and (ii) $E_{i}=12.5$ eV, located in area B.
@@
For the area A, where SOA and ExA have the same sign but different magnitudes, we used $E_{i}=$10 eV as a representative energy (marked as A in Fig. 5). This is the same incident energy used to record the energy-loss spectra presented in Fig. 3. The evolution of both SOA and ExA with $|Q|$ is presented in Fig. 6(a). The large variations of SOA and ExA indicate the strong competition between the two scattering processes responsible for these asymmetries. In particular, in the vicinity of $|Q|=0.4$-1 SOA is rather large and is comparable to ExA. Hence, it can influence the magnon excitation process. For one direction of magnetization, i.e., $[\overline{1}10]$ it would lead to a reduction of the magnon amplitude and for the opposite magnetization direction, i.e., $[1\overline{1}0]$ it would amplify the magnon amplitude. This is similar to the effect observed for the Co double layer on Ir(001) [15]. The main difference is that the sign of ExA is positive in the present case.

### v2

@@ Abstract
The propagation of magnons along a symmetry path may depend on the direction of propagation, similar to many other quasiparticles in nature. This phenomenon is commonly referred to as nonreciprocity. In addition to the fact that it is of great interest to understand the fundamental physical mechanism leading to this nonreciprocal propagation, the phenomenon of magnon nonreciprocity may be used to design magnon-based logic devices. Recently, it has been demonstrated that a significantly large spin-orbit coupling can lead to giant nonreciprocity of exchange-dominated terahertz magnons, when they are excited by means of spin-polarized electrons [Phys. Rev. Lett. 132, 126702 (2024)]. Here, by providing experimental results on two additional systems we demonstrate the generality of the observed phenomenon. Comparing the results of a Co/Ni bilayer on Ir(001) to those of a Co double layer on Ir(001) and W(110) we unravel the impact of the interfacial electronic hybridization on the observed phenomenon and provide further insights into the microscopic mechanism leading to this nonreciprocal magnon excitation. It was observed that the interfacial electronic hybridization is of some importance but is not crucial for the magnon nonreciprocity. This is an important observation since the electronic hybridization is known to be a key aspect in the determination of the magnetic properties at the interface. On the other hand the choice of the incident energy of the incoming electron beam is decisive for the observation of the effect. Our results indicate that depending on the energy of the incident electron beam and the scattering geometry the magnon nonreciprocity can be tuned and even be inverted for some ranges of the magnon momentum. The results may inspire new ideas for the generation of nonreciprocal ultrafast terahertz magnons in ultrathin films of $3d$ magnetic metals as well as in two-dimensional van der Waals magnets, interfaced with solids possessing a large spin-orbit coupling, for applications in the field of magnonics.
@@ I Introduction
Usually when a quasiparticle is propagating along two opposite but equivalent directions it is expected to exhibit identical properties, e.g., amplitude, lifetime, propagation speed, etc. [1, 2]. However, under some circumstances, e.g., in the presence of a chiral interaction, these properties may depend on the direction of propagation. This phenomenon is usually referred to as nonreciprocity and has been expected to be observed for magnons –the quanta of spin waves in magnetic materials– when they are propagating along two opposite directions of a symmetry path. The effect is of great importance from both fundamental as well as application points of view. From a fundamental point of view, a nonreciprocal propagation of a quasiparticle is associated with the presence of an interaction, which possibly breaks the time-reversal symmetry. It is, therefore, interesting to unravel the nature of the nonreciprocity, in order to address the presence of such a time-reversal symmetry breaking interaction. From an application point of view, nonreciprocal propagation of magnons is of great importance to realize novel magnon-based logic devices useful for applications in the field of magnonics and spintronics, as has already been demonstrated in the case of magnons with gigahertz frequencies [3, 4, 5, 6, 7, 8, 9, 10].
From the band magnetism point of view magnons in itinerant electron systems can be described by the coherent superposition of all possible electron-hole pairs for which the electron is of minority spin character and the hole is of majority character, formed across the Fermi-level. The wavefunction of such an excitation can be written as a linear superposition of the wavefunctions of all possible correlated electron-hole pairs, which satisfy this condition (the fact that the total angular momentum of a magnon is $1\hbar$), as well as the energy and wavevector conservation rules [11, 12]. It is straightforward to imagine that if the degeneracy of the bands is broken solely due to the exchange splitting, one would not expect a difference in the amplitude of the magnons’ wavefunction, in particular in the case of exchange-dominated magnons, with the opposite wavevector (or the two opposite directions of the magnetization, if the sample is ferromagnetic). The situation changes if one adds the relativistic spin-orbit coupling (SOC) into the formalism. In this case in addition to the exchange splitting one needs to consider the spin-orbit induced band splitting. Hence, one would expect a difference in the amplitude of the magnons’ wavefunction with the opposite signs of the wavevector (or for the two opposite directions of the magnetization). This is a direct consequence of the fact that the electrons (and holes) belong to spin-orbit spin-mixed bands now. In an extreme case scenario e.g., in the case of topological insulators, it has been predicted that due to the spin-momentum locking of the electronic states—a unique intrinsic feature of strongly spin-orbit coupled states in solids— even the collective charge excitations shall carry a spin character [13, 14].
From a broader perspective, one may consider a magnon as a bosonic quasiparticle which possesses three characteristics, i. e., (i) an eigenfrequency, (ii) a lifetime and (iii) an amplitude. In principle, SOC can lead to a magnon nonreciprocity in different ways.
First, SOC as a chiral interaction can break the energy degeneracy of the magnons having the same wavevector but propagate along opposite (but crystallographically equivalent) directions and, thereby, can lead to nonreciprocal magnons. However, this interaction does not directly act on magnons as bosons. We have demonstrated earlier that SOC leads to the presence of the antisymmetric Dzyaloshinskii–Moriya interaction (DMI), which can then influence the magnon eigenfrequency [15]. In a simplified picture SOC appears in the spin Hamiltonian in the form of DMI [15, 16, 17, 18].
Second, SOC can influence the magnon lifetime by modifying the Landau damping of the magnons having opposite propagation directions. In metallic magnets the lifetime of high-energy exchange-dominated magnons is governed by their dissipation into single-particle electron-hole pairs (also called Stoner pairs), known as Landau damping [19, 20, 21, 22]. Since in ultrathin ferromagnetic films grown on heavy-element metallic substrates SOC inside the ferromagnetic film can be large, the dissipation takes place within the spin-orbit spin-mixed electronic bands. It has been shown that this can lead to an asymmetric magnon lifetime [23]. SOC leads to an additional decay channel, which depends on the propagation direction of the magnons [23, 17].
Third, SOC should, in principle, influence also the amplitude of the excited magnons via the excitation process, depending on the sign of their wavevector. This phenomenon has just been demonstrated very recently [24]. We have shown that if the electrons are used as the excitation and probe particles, the phenomenon can be easily observed. The chiral SOC leads to a nonreciprocal excitation cross-section and, hence, nonreciprocal magnon amplitudes. This is due to the fact that SOC is a chiral interaction and in ultrathin ferromagnets the time-reversal $\mathcal{T}$ and inversion symmetry is broken. It is important to notice that such a nonreciprocity is caused by the fact that the excitation cross-section of magnons is not the same for the two opposite propagation directions (or the two opposite directions of magnetization). The effect is intimately related to the exchange process as the microscopic physical mechanism leading to the excitation of these exchange-dominated magnons. Hence, this kind of nonreciprocity is expected to be observed in all excitation schemes, which involve the exchange mechanism as the underlying mechanism for magnon excitation.
It is worth emphasizing that the magnon nonreciprocity is expected to be observed for the exchange-dominated terahertz (THz) magnons. Owing to the ultrafast propagation of these type of magnons, they would allow for information processing on ultrashort time scales [25]. Hence, such magnons are of great interest to the field of ultrafast magnonics and spintronics [26].
@@
In order to unravel the microscopic origin of the spin-orbit induced magnon nonreciprocity we investigated the impact of two important factors on the magnon nonreciprocity: (i) the interfacial electronic hybridization, and (ii) the atomic structure and the symmetry of the magnetic overlayer as well as the choice of the substrate.
In ferromagnetic films grown on heavy-element substrates, which are the main focus of the present work, it is generally accepted that SOC is mainly an interfacial effect and originates from the underlying heavy-element metallic substrates. In this respect, it is important to know what is the role of the interfacial electronic hybridization in the observed effect. Note that the role of the interfacial electronic hybridization is known to be extremely important in the determination of the magnetic properties of such ultrathin films. For instance, the interfacial magnetic moments and the Heisenberg exchange parameters depend strongly on the degree of the interfacial electronic hybridization [32]. In connection to this point we have recently investigated the role of the interfacial electronic hybridization on the strength and chirality of DMI [33]. Comparing the experimental results to those of first principles calculations it was observed that such a hybridization is crucial for DMI (both its strength and chirality) [33].
@@
Although the spin-orbit induced magnon nonreciprocity is expected to be observed by other excitation schemes, it should be best investigated by means of spin-polarized electron scattering experiments, e.g., spin-polarized high-resolution electron energy-loss spectroscopy (SPHREELS). While the original experiments were performed on a Co double layer on Ir(001), in order to investigate the impact of the interfacial electronic hybridization on the magnon nonreciprocity, we replaced the interface Co atomic layer with a Ni layer. We note that both Co double layer and Co/Ni bilayer  [^1] The term double layer refers to the case in which the sample is composed of two atomic layers of the same material. Likewise, the term bilayer refers to the case in which the sample is composed of two atomic layers made of two different materials. possess nearly the same geometrical structure (crystallographic structure and the interatomic distances) [35, 36, 37]. Since most of the $3d$ states of Ni are occupied and are located below the Fermi level, it is expected that these states hybridize rather weakly with that of the $5d$ states of the Ir substrate [38, 33]. This would provide a unique possibility to examine the influence of the interfacial electronic hybridization on the observed magnon nonreciprocity. On the other hand considering the available $5d$ states of the substrate, the tungsten exhibits a lower number of occupied $5d$ states compared to iridium. Hence, it would be interesting to replace the Ir(001) substrate with W(110) and see how it changes the magnon nonreciprocity. Unlike Ir, which possesses a face-centered cubic structure, the crystal structure of W is body-centered cubic. Since in the case of Co double layer on W(110) the atomic arrangement is also different, one would, therefore, expect to observe effects associated with the change in the atomic arrangement as well. In order to address the above mentioned points we investigated the details of the spin-dependent electron scattering from two systems, i.e., Co/Ni/Ir(001) and 2Co/W(110) and compare the results to those of 2Co/Ir(001), investigated previously [24]. We will show that, although the interfacial electronic hybridization is crucial for the determination the magnetic properties of the magnetic layers grown on a nonmagnetic substrate, they are not decisive for the observed magnon nonreciprocity. In contrary, the magnon nonreciprocity depends strongly on the incident electron energy and the scattering geometry. Moreover, for some ranges of magnon wavevectors the magnon nonreciprocity can even be inverted.
@@
The paper is organized as following. In Sec. II we provide the experimental details. The substrate preparation, film growth and the SPHREELS experiments are explained. The results of the two systems, namely Co/Ni/Ir(001) and 2Co/W(110) are presented in Secs. III.1 and III.2, respectively, and are discussed in detail. A concluding remark with the prospective of the observation of the effect in other layered magnetic structures is provides in Sec. IV.
@@ II.1 Substrate preparation
In this study we used Ir(001) and W(110) as substrates. The choice of Ir(001) is based on the fact that one can grow well-ordered ultrathin films of $3d$ ferromagnetic metals on this substrate. In particular considering Co and Ni as overlayers the growth mode, lattice structure and the interatomic distances are very similar [35, 39, 40]. They both grow in a face-centered tetragonal structure [35, 39, 40]. This is particularly important to address the effects associated with the changes in the chemical properties of the interface layer without changing its geometrical structures. The choice of W(110) is based on the fact that it also exhibits a rather large SOC. Moreover, one can grow high-quality ultrathin films of $3d$ ferromagnets on this surface [41, 42, 43, 44, 45].
@@
Figure 1: The magneto-optical Kerr effect (MOKE) hysteresis loops recorded on Co/Ni/Ir(001) [(a) and (b)] and 2Co/W(110) [(c) and (d)] samples. The loops were recorded in the longitudinal geometry at room temperature. In both cases the external magnetic field was applied along the [$\bar{1}$0]-direction of the substrates. The red and blue colors indicate the upward and downward magnetic field sweep, respectively. Panels (b) and (d) represent a magnified part of panels (a) and (c), respectively.
@@ II.2 Film deposition and characterization
Prior to the spectroscopy experiments the sample was fully magnetized using a magnetic field of about $2.5$ kOe. The reason for such a high magnetic field is to avoid the presence of possible magnetic domains in the sample. Note that during the spectroscopy experiments no magnetic field was applied. It is, therefore, crucial to ensure that the sample is in the single domain state.
@@ II.3 Spin-polarized high-resolution electron energy-loss spectroscopy
It is well-known that the magnon excitation process in SPHREELS is based on the exchange process [53, 54, 55, 49]. Since a magnon carries a total angular momentum of $1\hbar$ it can only be excited by incidence of a minority electron (when $\bm{P}$ is parallel to $\bm{M}$). This means that in fact the difference spectra $I_{\mu\downarrow}-I_{\mu\uparrow}$, which describe a spin-flip process, shall contain all the information regarding the magnons as spin-flip excitations. Since the magnon excitation is mediated by the exchange process, in the absence of SOC the excitation amplitude should be independent from the direction of $\bm{M}$ with respect to the crystallographic directions. It should only depend on the relative orientation of $\bm{P}$ and $\bm{M}$. Hence, the intensity of the difference spectra is expected to be equal for the two magnetization directions, despite a negative sign. The presence of SOC during the scattering event leads to an inequality of the intensity of the difference spectra, since it couples the spin to the crystallographic directions. On the other hand reversing the direction of $\bm{M}$ is equivalent to reversing the direction of the magnon wavevector [15, 23, 45]. This is because in a ferromagnet the time-reversal symmetry $\mathcal{T}$ is naturally broken. Hence, reversing the direction of $\bm{M}$ is equivalent to a time-reversal experiment. Therefore, the magnons for two opposite magnetization directions shall represent cases in which the magnons propagate along opposite directions. In principle, in the experiment changing the sign (or direction) of $\bm{Q}$ can be accomplished in two ways: (i) by changing the scattering geometry and (ii) by reversing the direction of $\bm{M}$ [15, 23]. Since changing the scattering geometry would change the scattering matrix elements, this may lead to unwanted effects. In order to avoid such effects one can simply reverse the direction of $\bm{M}$, keeping the scattering geometry unchanged (see Fig. 2). It has been shown that the presence of SOC leads to the fact that the magnon excitation amplitude, as seen by SPHREELS, is no longer equal for the two opposite directions of $\bm{M}$. The phenomenon is understood in terms of a strong competition between spin-orbit and exchange scattering mechanisms [24]. While in most of ferromagnetic elements such as Fe, Co and Ni SOC is not significantly large, the spin-orbit scattering is expected to be small. However, in ultrathin films of ferromagnetic metals grown on substrates with a large SOC spin-orbit scattering can be significant. The presence of a large SOC can lead to an inequality of magnon amplitudes, a phenomenon known as magnon nonreciprocity. In order to understand the effect in detail we prepared and investigated two additional systems as will be discussed in the following section.
@@ III.1 The case of Co/Ni/Ir(001)
The observed magnon nonreciprocity is also present for magnons having other wavevectors (see the discussion below). The difference spectra presented in Fig. 3 contain all the information regarding the excited magnon wavepackets, e.g., amplitude, lifetime, frequency, group and phase velocity, etc. One may, therefore, construct the magnon wavepackets propagating along two opposite directions based on these information [56, 20]. Since the excitation energy (the magnon eigenfrequency), the lifetime, group and phase velocities are nearly the same for the two wavepackets, we realized that the main factor, which leads to the nonreciprocal propagation of the magnon wavepackets is due to the inequality of the amplitude of the wavepackets in the initial stage of propagation (just at the time when they are excited). It is important to notice that the nonreciprocal behavior of magnons can be an overall effect of the nonreciprocity in all characteristics of magnons, i.e., amplitude, frequency and lifetime. However, in the present case the nonreciprocity is mainly due to the nonreciprocal magnon amplitude. The nonreciprocity in the magnon frequency is rather small (on the order of 1 meV or less for this wavevector). This is due to the presence of the so-called chirality-inverted DMI. The effect has been discussed in detail in Refs. [37, 33]. The amplitude of the magnon wavepackets is directly given by the height of the difference spectra at the magnon excitation energy. Hence, it would be very useful to analyze, in detail, the dependence of the ratio of the peak heights of magnon excitation spectra for $\bm{M}\parallel[1\bar{1}0]$ and $\bm{M}\parallel[\bar{1}10]$ [see Fig. 3(c)]. This quantity may also be called as the magnon nonreciprocity. For a reciprocal propagation one expects a value equal to unity and for a unidirectional propagation this value should be zero. Since the absolute value of the intensity of the excitation spectra depends on the scattering geometry one should use the normalized difference spectra [similar to those shown in Fig. 3(c)]. In Fig. 4 the excitation peak intensities for the two magnetization directions as well as their ratio are presented for different values of $Q$. One immediate conclusion from the data presented in Fig. 4 is that the magnon nonreciprocity depends strongly on $Q$ and can even be inverted for some ranges of $Q$. We may emphasize that the definition of the nonreciprocity is independent of the excitation method. It is valid for all excitation schemes, which involve the fundamental exchange process as the microscopic physical mechanism behind the magnon generation.
@@
Figure 5: Spin asymmetries of elastically scattered electrons against the incident electron energy $E_{i}$. $A_{\bm{M}\parallel[1\bar{1}0]}$ and $A_{\bm{M}\parallel[\bar{1}10]}$ were recorded for two opposite directions of magnetization, $[1\bar{1}0]$ and $[\bar{1}10]$, respectively, as described in the text. Different contributions to the spin asymmetries are $A_{SO}$ and $A_{Ex}$ and are given by Eqs. (1) and (2), respectively. The gray solid curve represents the spin-orbit asymmetry recorded on the bare Ir(001) surface. Note that the scale of the right axis is by a factor of 2 larger, compared to the left axis.
@@
Data shown in Fig. 5 indicate that both SOA and ExA exhibit a strong dependence on $E_{i}$. The behavior is very similar to that of the Co double layer on the same substrate [24]. In particular, the pattern of SOA is very similar in both systems, indicating that the SOC is mainly supported by the Ir(001) substrate. The $E_{i}$-dependence of SOA recorded on the bare Ir(001) substrate is also shown for a comparison (see the right axis of Fig. 5). It is important to notice that although the intrinsic SOC of the magnetic layers is not negligible and might be of some importance, it should be substantially enhanced in the presence of Ir as a heavy metal. Since the interfacial electronic hybridization is expected to be marginal in the case of Ni/Ir interface, another important conclusion from this observation is that such a hybridization does not influence the behavior of SOA, drastically. In order to carefully analyze the spin asymmetries one may consider several different regions of $E_{i}$. For instance, below 7.5 eV, $A_{SO}$ and $A_{Ex}$ possess a similar magnitude but an opposite sign (shown by shaded yellow color in Fig. 5). In the range of 7.5 eV$<E_{i}<11.5$ eV $A_{SO}$ and $A_{Ex}$ have the same sign but a different magnitude (light-blue shaded area in Fig. 5). For the incident energies above 11.5 eV $A_{SO}$ and $A_{Ex}$ exhibit again an opposite sign, similar to the first region (yellow shaded area in Fig. 5). In order to assess the consequences of the incident energy on the magnon nonreciprocity we analyzed the behavior of these asymmetries as a function of $Q$ for two areas marked by A and B in Fig. 5. Note that one could also analyze the data for $E_{i}<7.5$ eV. The results are expected to be similar to those of area B (see the data provided for the 2Co/W(110) system in Sec. III.2). Both SOA and ExA were measured as a function of $Q$ for two different values of the incident energy (i) $E_{i}=10$ eV, located in area A, and (ii) $E_{i}=12.5$ eV, located in area B. Note that one can, in principle, choose any point within these areas as a representative point.
@@
For the area A, where SOA and ExA have the same sign but different magnitudes, we used $E_{i}=$10 eV as a representative energy (marked as A in Fig. 5). This is the same incident energy used to record the energy-loss spectra presented in Fig. 3. The evolution of both SOA and ExA with $|Q|$ is presented in Fig. 6(a). The large variations of SOA and ExA indicate the strong competition between the two scattering processes responsible for these asymmetries. In particular, in the vicinity of $|Q|=0.4$-1 SOA is rather large and is comparable to ExA. Hence, it can influence the magnon excitation process. For one direction of magnetization, i.e., $[\overline{1}10]$ it would lead to a reduction of the magnon amplitude and for the opposite magnetization direction, i.e., $[1\overline{1}0]$ it would amplify the magnon amplitude. This is similar to the effect observed for the Co double layer on Ir(001) [24]. The main difference is that the sign of ExA is positive in the present case. Note that the choice of $|Q|=0.4$-1 is based on the fact that the magnon signal for this wavevector can easily be resolved. The effect is expected to be even larger at smaller magnon wavevectors.
@@
As a side remark, it is important to emphasize that the magnon nonreciprocity is not directly a consequence of DMI, since it is observed to be large in systems with a small DMI-induced magnon nonreciprocity in the frequency (for instance the Co/Ir and Ni/Ir interfaces, which exhibit a chirality-inverted DMI [52, 38, 33]). Moreover, the effect becomes more significant at lower wavevectors as seen in Figs. 4 and 6 (see also Sec. IV).